OpenStudy (anonymous):

a drawer contains 4 red socks ,3 white socks,and 3 blue socks . without looking , you select a sock at random, replace it,and select a second sock at random. what is the probability that the first sock is blue and the second sock is red?

OpenStudy (lgbasallote):

there are 10 socks in total the probability you pick a blue sock is 3/10 the probability you pick a red sock is 4/10 <--10 because you re-place the sock you took therefore it's \[\frac{3}{10} \times \frac{4}{10}\]

OpenStudy (lgbasallote):

you got the logic right?

OpenStudy (anonymous):

yea but do i add of mutiply the 3,4

OpenStudy (lgbasallote):

you multiply

OpenStudy (lgbasallote):

probability is usually multiplication i believe

OpenStudy (anonymous):

well that would put it at 12/100 and my choices are 3/5,7/20,3/25,or 7/100

OpenStudy (lgbasallote):

you can still simplify 12/100 ;)

OpenStudy (anonymous):

3/5

OpenStudy (lgbasallote):

try to check is 12 divided by 3 equal to 100 divided by 5?

OpenStudy (lgbasallote):

Hint!! \[\frac{4}{10} = \frac{2}{5}\]

OpenStudy (anonymous):

7/20

OpenStudy (lgbasallote):

you're guessing arent you

OpenStudy (lgbasallote):

lol

OpenStudy (anonymous):

yes sorry not understanding i hate when i don't get it

OpenStudy (lgbasallote):

look at \[\frac{12}{100}\] try to divide 12 AND 100 by 4 what do you get? \[\frac{12 \div 4}{100 \div 4}\]

OpenStudy (anonymous):

ok i see now thanks u are smart

OpenStudy (lgbasallote):

hehe ^_^ just keep on practicing ;)

OpenStudy (anonymous):

I have one more like that if u have a mintue could u walk me thourgh it step by step?

OpenStudy (lgbasallote):

well i have to go in a few minutes..i better not..i wouldnt want to leave in the middle lol

OpenStudy (anonymous):

ok thanks